# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ D.1 Useful Lemmas
$\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]\overset{(i)}{\leq}$ $\alpha^{2}\left(\sigma^{2}+\left\|\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)\right\|^{2}\right)-2\alpha\langle\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right),\theta_{k,t}-\theta_{*}\rangle+\left\|\theta_{k,t}-\theta_{*}\right\|^{2}$
where $(i)$ follows from the definition of the variance ${\rm Var[X]}=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$, $(ii)$ follows from the fact that the $\frac{b\mu}{\exp\beta}$-strong convexity and $\frac{BL}{\exp\beta}$-smoothness of $\mathcal{L}_{\rm FT}({\phi},{\theta})$ w.r.t. $\theta$ in Lemma 1 deduces that $\langle\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)-\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{*}\right),\theta_{k,t}-\theta_{*}\rangle\geq\frac{b\mu BL}{(b\mu+BL)\exp\beta}\|\theta_{k,t}-\theta_{*}\|^{2}+\frac{\exp\beta}{b\mu+BL}\|\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)\|^{2}$ (Theorem 2.1.12 in (Nesterov et al., 2018)), where $\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{*}\right)=0$. Using the condition $\alpha\leq\frac{2\exp\beta}{b\mu+BL}$ and unconditioning on $\theta_{k,t-1}$, we have Eq. (40). ∎
D.2 Proof of Theorem 1
Proof.
Let $\mathbf{L}(\phi_{k})=\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}[\mathcal{L}_{\rm PL}(\phi_{k},\theta^{T}_{*})]$ and let ~$\widetilde{\nabla}\mathbf{L}(\phi_{k})=\frac{\partial\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}[\mathcal{L}_{\rm PL}(\phi_{k},{\theta}_{k,D}^{T})]}{\partial\phi_{k}}$, ^$\widehat{\nabla}\mathbf{L}(\phi_{k})=\frac{\partial\ell_{\rm PL}(\phi_{k},{\theta}_{k,D}^{T};z_{k})}{\partial\phi_{k}}$, then using the smoothness assumption of , we have
@@ D.2 Proof of Theorem 1
${\bf L}(\phi_{k+1})\leq$ ${\bf L}(\phi_{k})+\left\langle\nabla{\bf L}(\phi_{k}),\phi_{k+1}-\phi_{k}\right\rangle+\frac{P}{2}\|\phi_{k+1}-\phi_{k}\|^{2}$
@@
$\leq$ ${\bf L}(\phi_{k})-\eta\left\langle\nabla{\bf L}(\phi_{k}),\widehat{\nabla}{\bf L}(\phi_{k})-\nabla{\bf L}(\phi_{k})\right\rangle-\eta\left\|{\nabla}{\bf L}(\phi_{k})\right\|^{2}+\eta^{2}P\|\nabla{\bf L}(\phi_{k})\|^{2}+\eta^{2}P\|\nabla{\bf L}(\phi_{k})-\widehat{\nabla}{\bf L}(\phi_{k})\|^{2}$
$\leq$ ${\bf L}(\phi_{k})-\left(\frac{\eta}{2}-\eta^{2}P\right)\|\nabla{\bf L}(\phi_{k})\|^{2}+\left(\frac{\eta}{2}+\eta^{2}P\right)\|{\nabla}{\bf L}(\phi_{k})-\widehat{\nabla}{\bf L}(\phi_{k})\|^{2}$
$\overset{(i)}{\leq}$ ${\bf L}(\phi_{k})-\left(\frac{\eta}{2}-\eta^{2}P\right)\|\nabla{\bf L}(\phi_{k})\|^{2}+\left(\frac{\eta}{2}+\eta^{2}P\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right],$
where $(i)$ follows from Jensen’s inequality. Taking expectation on both sides and rearranging the terms conditioning on $\eta(\frac{1}{2}-\eta P)\geq 0$, we have
$\left(\frac{\eta}{2}-\eta^{2}P\right)\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]$
$\leq$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right]$
$\leq$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},\theta_{k,D}^{T})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right]$
$+\left(\frac{\eta}{2}+\eta^{2}P\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},\theta_{k,D}^{T})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}\right\|^{2}\right]$
$\overset{(ii)}{\leq}$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P\right)\sigma_{3}+\frac{8M^{2}T^{2}}{\beta^{2}}\left(\frac{\eta}{2}+\eta^{2}P\right)\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)$
$+8M^{2}\left(\frac{\eta}{2}+\eta^{2}P\right)\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t},$
@@
$\left(\frac{\eta}{2}-\eta^{2}P\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]$
$\leq$ $\frac{{\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)}{K}+\left(\frac{\eta}{2}+\eta^{2}P\right)\sigma_{3}^{2}+\frac{8M^{2}T^{2}}{\beta^{2}}\left(\frac{\eta}{2}+\eta^{2}P\right)\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)$
$+8M^{2}\left(\frac{\eta}{2}+\eta^{2}P\right)\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}$
$\overset{(i)}{\leq}$ $\frac{{\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)}{K}+\left(\frac{\eta}{2}+\eta^{2}P\right)\sigma_{3}^{2}+\frac{8M^{2}T^{2}}{\beta^{2}}\left(\frac{\eta}{2}+\eta^{2}P\right)+\frac{8M^{2}}{\kappa}\left(\frac{\eta}{2}+\eta^{2}P\right)\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)$
@@
where $(i)$ follows from that $D\geq-\frac{\log\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)}{\log\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)}$ implies $\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\leq 1$, $\beta\leq\log\left(\alpha b\mu/1-\sqrt{\frac{1-\kappa}{3}}\right)$ implies $\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\leq\frac{1}{1-3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}}\leq\frac{1}{\kappa}$, for some positive constant $\kappa<1$
Then, setting $\eta=\frac{1}{4P}$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]\leq$ $\frac{16P\left({\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)\right)}{K}+3\sigma_{3}^{2}+\frac{24M^{2}T^{2}}{\beta^{2}}$
$+\frac{24M^{2}}{\kappa}\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)$
@@
Then, denoting $\tau=3\sigma_{3}^{3}+\frac{144M^{2}\alpha^{2}\sigma_{1}^{2}}{\kappa}$, we obtain the result of Theorem 1. ∎
Appendix E Further Details on the Experimental Setup
E.1 GPT-4 prompts for generating preference data for sentiment generation

### v2

@@ 0.D.1 Useful Lemmas
$\overset{(ii)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\left(\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]\right)$
$+\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right),$
where $(i)$ follows from Lemma 6, $(ii)$ follows from a constrain on $\beta$ such that $\beta\leq\log\frac{4\alpha B^{2}(L^{2}\mu^{2}+M^{2}\rho^{2})}{b\mu(2L+M^{2})^{2}}$.
Telescoping over $t$ from 00 to $D-1$, we have
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]\\ {\leq}&\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\left\|\frac{\partial\theta_{0,k}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\left\|\theta_{k,0}-\theta_{*}\right\|^{2}\right)\\ &+\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (41)
Applying the initialization $\theta_{k,0}=\theta_{0}$ at each inner-loop updating interation, we have the result in Eq. (38).
Lemma 8
Based on Assumption 1, 2 and 3, we have for any $k\in\{0,1,2,\ldots,K-1\}$,
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})}{\partial\phi_{k}}\right\|^{2}\right]\\ \leq&\frac{2M^{2}T^{2}(2L+M^{2})^{2}}{\beta^{2}\mu^{2}}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\\ &+8M^{2}\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (42)
Proof
@@
$\overset{(i)}{\leq}$ $2M\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|+\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|\left\|\nabla_{{\theta}}\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})-\nabla_{{\theta}}\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})\right\|$
$\overset{(ii)}{\leq}$ $2M\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|+\frac{MT(2L+M^{2})}{\beta\mu}\|\theta_{k,D}-{\theta_{*}}\|,$
where $(i)$ follows from Jensen’s inequality and Lemma 3, $(ii)$ follows from Jensen’s inequality and Lemma 3, 4.
Using the fact $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and taking expectation on both sides, we have
$\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})}{\partial\phi_{k}}\right\|^{2}\right]$
$\leq$ $8M^{2}\left(\mathbb{E}\left[\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\|\theta_{k,D}-{\theta_{*}}\|^{2}\right]\right),$
and using Lemma 7, we have Eq. (42).
0.D.2 Proof of Theorem 3.1
Proof
Let $\mathbf{L}(\phi_{k})=\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}[\mathcal{L}_{\rm PL}(\phi_{k},\theta^{T}_{*})]$ and let ~$\widetilde{\nabla}\mathbf{L}(\phi_{k})=\frac{\partial\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}[\mathcal{L}_{\rm PL}(\phi_{k},{\theta}_{k,D}^{T})]}{\partial\phi_{k}}$, ^$\widehat{\nabla}\mathbf{L}(\phi_{k})=\frac{\partial\ell_{\rm PL}(\phi_{k},{\theta}_{k,D}^{T};z_{k})}{\partial\phi_{k}}$. Adopting Lemma 5 with Jensen’s inequality, we have the $P_{\phi}$-smoothness of $\mathbf{L}(\phi_{k})$, and thus
${\bf L}(\phi_{k+1})$
$\leq$ ${\bf L}(\phi_{k})+\left\langle\nabla{\bf L}(\phi_{k}),\phi_{k+1}-\phi_{k}\right\rangle+\frac{P_{\phi}}{2}\|\phi_{k+1}-\phi_{k}\|^{2}$
@@
$\leq$ ${\bf L}(\phi_{k})-\eta\left\langle\nabla{\bf L}(\phi_{k}),\widehat{\nabla}{\bf L}(\phi_{k})-\nabla{\bf L}(\phi_{k})\right\rangle-\eta\left\|{\nabla}{\bf L}(\phi_{k})\right\|^{2}$
$+\eta^{2}P\|\nabla{\bf L}(\phi_{k})\|^{2}+\eta^{2}P_{\phi}\|\nabla{\bf L}(\phi_{k})-\widehat{\nabla}{\bf L}(\phi_{k})\|^{2}$
$\leq$ ${\bf L}(\phi_{k})-\left(\frac{\eta}{2}-\eta^{2}P_{\phi}\right)\|\nabla{\bf L}(\phi_{k})\|^{2}+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\|{\nabla}{\bf L}(\phi_{k})-\widehat{\nabla}{\bf L}(\phi_{k})\|^{2}$
$\overset{(i)}{\leq}$ ${\bf L}(\phi_{k})-\left(\frac{\eta}{2}-\eta^{2}P_{\phi}\right)\|\nabla{\bf L}(\phi_{k})\|^{2}+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right],$
where $(i)$ follows from Jensen’s inequality. Taking expectation on both sides and rearranging the terms conditioning on $\eta(\frac{1}{2}-\eta P_{\phi})\geq 0$, we have
$\left(\frac{\eta}{2}-\eta^{2}P_{\phi}\right)\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]$
$\leq$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right]$
$\leq$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},\theta_{k,D}^{T})}{\partial\phi_{k}}-\widehat{\nabla}\mathbf{L}(\phi_{k})\right\|^{2}\right]$
$+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\mathbb{E}_{T\sim\mathcal{D}_{\mathcal{T}}}\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},\theta_{k,D}^{T})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}(\phi_{k},{\theta}^{T}_{*})}{\partial\phi_{k}}\right\|^{2}\right]$
$\overset{(ii)}{\leq}$ $\mathbb{E}\left[{\bf L}(\phi_{k})\right]-\mathbb{E}\left[{\bf L}(\phi_{k+1})\right]+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\sigma_{3}$
$+\!\!8M^{2}\!\!\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\!\!\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\!\!+\!\!6\alpha^{2}\!\!\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\!\!\right)\!\!\sum_{t=0}^{D-1}\!\!\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}$
$+\frac{2M^{2}T^{2}(2L+M^{2})^{2}}{\beta^{2}\mu^{2}}\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right),$
@@
$\left(\frac{\eta}{2}-\eta^{2}P_{\phi}\right)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]$
$\leq$ $\frac{{\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)}{K}+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\sigma_{3}^{2}$
$+\!\!8M^{2}\!\!\left(\frac{\eta}{2}\!\!+\!\!\eta^{2}P_{\phi}\right)\!\!\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\!\!+\!\!6\alpha^{2}\!\!\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\!\!\right)\!\!\sum_{t=0}^{D-1}\!\!\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}$
$+\frac{2M^{2}T^{2}(2L+M^{2})^{2}}{\beta^{2}\mu^{2}}\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)$
$\overset{(i)}{\leq}$ $\frac{{\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)}{K}+\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\sigma_{3}^{2}$
$+\!\!\frac{8M^{2}}{\kappa}\!\!\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right)\!\!\left(\!\!\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\!\!+\!\!6\alpha^{2}\!\!\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\!\!\right)$
$+\frac{2M^{2}T^{2}(2L+M^{2})^{2}}{\beta^{2}\mu^{2}}\left(\frac{\eta}{2}+\eta^{2}P_{\phi}\right),$
@@
where $(i)$ follows from that $D\geq-\frac{\log\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)}{\log\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)}$ implies $\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\!\!\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\leq 1$, $\beta\leq\log\left(\alpha b\mu/1-\sqrt{\frac{1-\kappa}{3}}\right)$ implies $\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\leq\frac{1}{1-3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}}\leq\frac{1}{\kappa}$, for some positive constant $\kappa<1$.
Then, setting $\eta=\frac{1}{4P_{\phi}}$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla{\bf L}(\phi_{k})\|^{2}\right]$
$\leq$ $\frac{16P_{\phi}\left({\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)\right)}{K}+3\sigma_{3}^{2}+\frac{6M^{2}T^{2}(2L+M^{2})^{2}}{\beta^{2}\mu^{2}}$
$+\frac{24M^{2}}{\kappa}\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\right)$
$=$ $\frac{16P_{\phi}\left({\bf L}(\phi_{0})-\inf_{\phi}{\bf L}(\phi)\right)}{K}+\left((2L+M^{2})^{2}+\frac{4\alpha^{2}\sigma^{2}(2L+M^{2})^{2}}{\kappa}+\frac{24M^{2}\sigma_{2}^{2}}{\kappa}\right)\frac{6M^{2}T^{2}}{\beta^{2}\mu^{2}}$
$+3\sigma_{3}^{3}+\frac{144M^{2}\alpha^{2}\sigma_{1}^{2}}{\kappa}$
@@
Then, denoting $\tau=3\sigma_{3}^{3}+\frac{144M^{2}\alpha^{2}\sigma_{1}^{2}}{\kappa}$, we obtain the result of Theorem 3.1.
Appendix 0.E Further Details on the Experimental Setup
0.E.1 Statistics of datasets
Tab. 4 lists statistics of the used training and evaluation data for CSG and KAG across different domains.
Table 4: Dataset statistics.
Domain Training Data Evaluation Data
Fine-Tuning Pref. Learn. Reference Preference
CSG
book 1,000 500 500 500
dvd 1,000 500 500 500
electronics 1,000 500 500 500
kitchen 1,000 500 500 500
KAG
askacademia (ac) 31,450 2,095 1,708 1,708
askanthropology (an) 3,910 203 268 268
askbaking (ba) 44,007 2,096 268 268
askcarguys (ca) 3,227 159 117 117
askculinary (cu) 45,710 2,094 2,563 2,563
askdocs (do) 6,449 315 455 455
askengineers (en) 57,096 3,154 2,638 2,638
askhistorians (hi) 3,264 113 164 164
askhr (hr) 8,295 641 395 395
askphilosophy (phi) 10,307 608 677 677
askphysics (phy) 7,364 409 587 587
askscience (sci) 13,316 899 977 977
asksciencefiction (scif) 29,382 1,576 1,987 1,987
asksocialscience (so) 2,706 147 188 188
askvet (ve) 3,300 170 224 224
changemyview (ch) 38,173 1,637 1,836 1,836
explainlikeimfive (ex) 19,592 1,014 1,070 1,070
legaladvice (le) 21,170 1,106 1,011 1,011
0.E.2 GPT-4 prompts for generating preference data for sentiment generation